Ten years ago, a father was 8 times the age of the son. if the total of their ages is now 56, what are their present ages?
Answer by lwsshak3(11628)

Ten years ago, a father was 8 times the age of the son. if the total of their ages is now 56, what are their present ages?
***
let x=present age of father
56-x=present age of son
10 yrs ago:
x-10=age of father
56-x-10=46-x=age of son
age of father=8 times age of son
x-10=8(46-x)
x-10=368-8x
9x=378
x=42
56-x=14
present age of father=42
present age of son=14